You make the following five end-of-year deposits into your investment account that earns 9% annual interest: $600 in year 1, $700 in year 2, $500 in year 3, $300 in year 4, and $400 in year 5. Note that the first cash flow of $600 which occurs at the end of year 1 earns interest for four years.What is the balance in your investment account at the end of year 5? Round to the nearest dollar amount.
Answer #1
Year Deposit Formula Future value
1 600 =600*(1+9%)^(5-1)                   847
2 700 =700*(1+9%)^(4-1)                   907
3 500 =500*(1+9%)^(3-1)                   594
4 300 =300*(1+9%)^(2-1)                   327
5 400 =400*(1+9%)^(1-1)                   400
Total                3,075
Total Investment at the end of year 5=$3,075